# A Survey on Large Language Model Acceleration based on KV Cache Management

### v2

@@ Abstract
Large Language Models (LLMs) have revolutionized a wide range of domains such as natural language processing, computer vision, and multi-modal tasks due to their ability to comprehend context and perform logical reasoning. However, the computational and memory demands of LLMs, particularly during inference, pose significant challenges when scaling them to real-world, long-context, and real-time applications. Key-Value (KV) cache management has emerged as a critical optimization technique for accelerating LLM inference by reducing redundant computations and improving memory utilization. This survey provides a comprehensive overview of KV cache management strategies for LLM acceleration, categorizing them into token-level, model-level, and system-level optimizations. Token-level strategies include KV cache selection, budget allocation, merging, quantization, and low-rank decomposition, while model-level optimizations focus on architectural innovations and attention mechanisms to enhance KV reuse. System-level approaches address memory management, scheduling, and hardware-aware designs to improve efficiency across diverse computing environments. Additionally, the survey provides an overview of both text and multimodal datasets and benchmarks used to evaluate these strategies. By presenting detailed taxonomies and comparative analyses, this work aims to offer useful insights for researchers and practitioners to support the development of efficient and scalable KV cache management techniques, contributing to the practical deployment of LLMs in real-world applications. The curated paper list for KV cache management is in: https://github.com/TreeAI-Lab/Awesome-KV-Cache-Management.
@@ 1 Introduction
Large Language Models (LLMs) [1, 2], trained on massive corpora, have revolutionized various domains such as natural language processing [3, 4, 5], computer vision [6, 7, 8], and multi-modal [9, 10, 11] tasks. Their ability to understand context and perform logical reasoning has enabled remarkable success in various fields, such as time series analysis [12, 13], recommendation [14, 15], autonomous driving [16, 17, 18], and healthcare [19, 20]. These breakthroughs are powered by state-of-the-art architectures and training paradigms, enabling models to achieve unparalleled performance across diverse tasks. Prominent LLMs, such as GPT [21, 22, 23], LLaMA [24, 25], DeepSeek [26, 27, 28], Mistral [29, 30], and GLM [31, 32], are built on the foundational transformer architecture [33], which excels at capturing long-range dependencies in sequential data. However, despite their powerful capabilities, the computational and memory demands of LLMs, particularly during inference, present significant challenges when scaling them to real-world, long-context, and real-time applications.
@@
A critical bottleneck in LLM inference lies in the efficient management of Key-Value (KV) pairs. Recently, caching techniques [34, 35] have been extensively employed to store previously computed intermediate results, allowing their reuse in subsequent inference steps to accelerate the model, such as graph neural networks [36, 37, 38]. Fortunately, the auto-regressive generation mechanism inherent to LLMs presents an opportunity to leverage KV caching for efficient text generation. Specifically, auto-regressive generation enables LLMs to produce text token by token, with each token conditioned on all previously generated ones. While this approach is highly effective for generating coherent and contextually relevant outputs, it suffers from poor scalability with long input sequences, as the computational and memory requirements grow quadratically with sequence length. The KV cache addresses this issue by storing key and value matrices from previous decoding steps, enabling their reuse and significantly reducing redundant computations.
@@
Several recent surveys [2, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50] have explored the domain of efficient LLMs. These surveys primarily examine various aspects of LLM efficiency, presenting valuable insights while leaving room for further refinement and innovation. In particular, many of these works primarily focus on holistic approaches to improving LLM efficiency, examining a wide range of techniques across multiple dimensions, such as span data-level optimizations (e.g., prompt engineering), model architecture-level optimizations (e.g., efficient transformer designs), and system-level optimizations (e.g., task scheduling). For instance, Ding et al.[42] explore efficiency techniques that integrate data-level and model architecture perspectives, while Miao et al.[43] examine efficient LLM inference from a comprehensive system-level perspective. Similarly, Tang et al.[46], Wan et al.[44], and Xu et al. [48] provide analyses that encompass data, model, and system-level optimizations, reflecting holistic approaches to LLM acceleration.
@@
On the other hand, some surveys focus on more specialized aspects for LLM acceleration. For example, Zhu et al.[2], Park et al.[40], Wang et al.[41], and Tang et al.[46] focus on model compression as a key aspect of model-level optimization. Similarly, Kachris et al.[47] examine hardware acceleration strategies tailored for LLMs, while Xu et al.[48] investigate parameter-efficient tuning approaches. Albalak et al.[49] discuss data selection strategies to enhance the efficiency of LLM training, and Xia et al.[51] highlight collaborative techniques, such as speculative decoding [52, 53], to accelerate model inference. Li et al. [54] focus on prompt compression. Similar to our work, Shi et al.[55], Li et al.[56], and Yuan et al. [57] also explore the use of KV caches to accelerate LLMs. However, our survey is both complementary and more comprehensive, offering a detailed taxonomy of KV cache management for text-based and multi-modal LLMs. We categorize techniques across token-level, model-level, and system-level perspectives and include benchmarks for both text and multi-modal scenarios. In particular, complementing existing KV cache surveys, we provide a detailed comparison of the differences and advantages of existing models at the token-level, model-level, and system-level.
@@ 2.1 Transformer Architecture
Transformers [33] have become the backbone of LLMs due to their ability to efficiently capture long-range dependencies sequential data, such as text. This capability makes them particularly well-suited for tasks like machine translation, text generation, and image captioning. The transformer architecture follows an encoder-decoder structure, where most LLMs utilize only the decoder component. We first introduce the core components of the Transformer decoder and then describe the critical auto-regressive generation mechanism. Particularly, we do not describe certain components in transformer, such as normalization, as they do not impact the understanding of KV cache management.
@@ 2.1.1 Transformer Decoder
As shown in Figure 1, a decoder-based transformer architecture is composed of multiple stacked Transformer blocks, each designed to process sequential data effectively. Typically, a Transformer block consists of two core components, i.e., a Multi-Head Self-Attention (MHSA) mechanism and a Feed Forward Network (FFN). These blocks are arranged sequentially, where the output of one block is passed as input to the next. This iterative design allows the model to refine its understanding of the input sequence progressively, making it highly effective for tasks such as text generation and language modeling.
Positional Encoding. Before the input sequence is processed by the Transformer blocks, it undergoes a preprocessing phase. First, a tokenizer processes the input sentence $X$ by splitting it into discrete units, such as words or subwords. The resulting sequence can be represented as ${X}=[x_{1},x_{2},\cdots,x_{|X|}]$. These tokens are then mapped to dense vector representations using an embedding layer, i.e., $\mathbf{X}=\mathbf{I}_{X}\mathbf{E}^{\top}$, where $\mathbf{I}_{X}\in\{0,1\}^{n\times d_{\text{vocab}}}$ represents the one-hot vector of tokenized input $X$, $\mathbf{E}\in\mathbb{R}^{d_{\text{vocab}}\times d_{x}}$ is the embedding matrix, and $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{|X|}]\in\mathbb{R}^{n\times d_{x}}$ is the resulting matrix of embedded token representations. Since the Transformer architecture does not inherently account for the order of tokens in a sequence, positional encodings are added to the token embeddings $\mathbf{X}$ to incorporate positional information. This can be expressed as $\mathbf{X}=\mathbf{X}+PE(X)$, where $PE(X)\in\mathbb{R}^{n\times d_{x}}$ represents a function [59, 60, 61] (e.g., sine and cosine-based positional encoding) that generates positional embeddings for the input ${X}$.
@@
where $\mathbf{W}_{Q_{i}}\in\mathbb{R}^{d_{x}\times d_{k}}$, $\mathbf{W}_{K_{i}}\in\mathbb{R}^{d_{x}\times d_{k}}$ and $\mathbf{W}_{V_{i}}\in\mathbb{R}^{d_{x}\times d_{v}}$ are the learned weight parameters. Then, the self-attention operation is applied to each triple $(\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i})$, and obtain the output of the $i$-th attention head $\mathbf{Z}_{i}$ as follows:
@@
where $\sqrt{d_{k}}$ is a scaling factor to ensure the numerical stability. To capture diverse relationships, multiple attention with $h$ heads are applied to $\mathbf{X}$ in parallel, and their outputs are concatenated with one transformation as follows:
@@
where Concat is concatenation operation and $\mathbf{W}_{O}\in\mathbb{R}^{d_{v}\times d_{o}}$ is the trainable parameters.
Following the self-attention mechanism, the output is passed through a Feed Forward Network (FFN). The FFN is a fully connected neural network that applies two linear transformations separated by a nonlinear activation function $\sigma(\cdot)$ (e.g, ReLU [62]) :
@@
where $\mathbf{W}_{1}\in\mathbb{R}^{d_{o}\times d_{1}}$ and $\mathbf{W}_{2}\in\mathbb{R}^{d_{1}\times d_{2}}$ are two parameters. Also, $\mathbf{b}_{1}\in\mathbb{R}^{d_{1}}$ and $\mathbf{b}_{2}\in\mathbb{R}^{d_{2}}$ are two bias vectors.
@@ 2.2.1 Auto-regressive Generation with KV Cache
Formally, at decoding step $t$, the new token embedding $\mathbf{x}_{t}$ is used to compute the query vector $\mathbf{q}^{t}_{i}$, key vector $\mathbf{k}^{t}_{i}$, and value vector $\mathbf{v}^{t}_{i}$ as follows:
@@
where $\mathbf{\hat{K}}_{i}^{t-1}\in\mathbb{R}^{t-1\times d_{k}}$ and $\mathbf{\hat{V}}_{i}^{t-1}\in\mathbb{R}^{t-1\times d_{v}}$ represent the cached key and value matrices of tokens in $X[1:t-1]$. These cached matrices are then used in the scaled dot-product attention computation for token $x_{t}$. The attention output $\mathbf{z}^{t}_{i}$ for the token $x_{t}$ at step $t$ is calculated as:
@@ 2.2.2 Time and Space Complexity Analysis
Given a transformer-based $L$-layer LLM with $h$ attention heads per layer and an input sequence of length $X=[x_{1},\cdots,x_{t}]$, we analyze the time saved and the space required to store cached KV pairs. For simplicity, we assume the Keys and Values of $t_{c}$ tokens are stored for all heads across all LLM layers.
Saved Time. For each token, the saved computation time comes from avoiding the repeated computation of Keys and Values in Equation (1), self-attention result in Equation (2), and linear transformation in Equation (3). We omit the time analyze on operations in transformer that do not affect the understanding of KV cache acceleration, such as layer norm and position encoding.
  • QKV Computation. The time of computing Queries, Keys and Values for each token in Equation (1) is $\triangle_{1}=O(2d_{x}d_{k}+d_{x}d_{v})$.
  • Self-attention Result. Additionally, computing each attention result $\mathbf{z}_{i}$ in Equation (2) takes $O(t(d_{k}+d_{v}))$.
  • Linear Transformation. To merge the $h$ attention results in Equation (3) the time is $\triangle_{2}=O(hd_{v}+d_{v}d_{o})$.
@@
Extra Space. Compared to computation without caching, additional space is required to store the cached KV pairs for $t_{c}$ tokens across $h$ attention heads and $L$ layers. Assuming each Key and Value is stored in Float16 precision, the total extra space needed can be expressed as:
$O(L\cdot h\cdot t_{c}\cdot 2\cdot sizeof(Float16))$ (11)
@@ 2.3 Challenges in KV Cache Management
  • Cache Eviction Policies: Determining which items to evict when the cache reaches its capacity is a complex problem. Popular policies [35] like Least Recently Used (LRU) or Least Frequently Used (LFU) do not align with LLMs patterns, leading to suboptimal performance.
@@ 3 Taxonomy
In the above sections, we analyzed how the number of cached Key-Value (KV) pairs significantly impacts both the computation time and the additional memory required during inference. Efficient KV cache management is critical to balancing performance improvements and resource utilization, especially as sequence lengths and model sizes continue to grow. After carefully reviewing existing approaches, we categorize KV cache optimization strategies into three levels: token-level optimization, model-level optimization, and system-level optimizations. Each level addresses specific aspects of the challenges associated with KV cache management and offers distinct techniques to enhance efficiency. The detailed taxonomy is illustrated in Fig. 2.
  • Token-Level Optimization refers to improving KV cache management efficiency by focusing on fine-grained the careful selection, organization, and compression at the token level, requiring no architectural changes to the original model. While KV cache selection (Sec. 4.1) focuses on prioritizing and storing only the most relevant tokens. KV cache budget allocation (Sec. 4.2) dynamically distributes memory resources across tokens to ensure efficient cache utilization under limited memory. Furthermore, KV cache merging (Sec. 4.3) reduces redundancy by combining similar or overlapping KV pairs, while KV Cache Quantization (Sec. 4.4) minimizes the memory footprint by reducing the precision of cached KV pairs. Finally, KV cache low-rank decomposition (Sec. 4.5) uses low-rank decomposition technique to reduce cache size.
  • Model-level Optimization refers to designing an efficient model structure to optimize KV cache management. This can further refer to several strategies: Attention grouping and sharing (Sec. 5.1) methods examine the redundant functionality of key and values and group and share KV cache within or across transformer layers. Architecture alterations (Sec. 5.2 emerge to design new attention mechanisms or construct extrinsic modules for KV optimization. Furthermore, there are also works designing or combining non-transformer architectures 5.3 that adopt other memory efficient designs like recurrent neural networks to optimize the KV cache in traditional transformers.
@@
Decomposition (Sec 4.5.2), dnode [ DecoQuant [65] , wnode_1 ] ] [Singular Value Decomposition (Sec 4.5.1), dnode [ ECKVH [66], EigenAttention [67], ZDC [68], LoRC [69], ShadowKV [70], Palu [71] , wnode_1 ] ] ] [KV Cache
@@
Merging (Sec 4.3.1), dnode [ CCM [101], LoMA [102], DMC [103], CaM [104], D2O [105], AIM [106], Look-M [107], KVMerger [108], CHAI [109], wnode_1 ] ] ] [KV Cache Budget Allocation (Sec. 4.2), dnode [Head-wise Budget Allocation (Sec 4.2.2), dnode [ AdaKV [110], CriticalKV [111], LeanKV [112], RazorAttention [113], HeadKV [114], DuoAttention [115] ,wnode_1 ] ] [Layer-wise Budget Allocation (Sec 4.2.1), dnode [ PyramidKV [116], PyramidInfer [117], DynamicKV [118], PrefixKV [119], SimLayerKV [120] , wnode_1 ] ] ] [KV Cache
Selection (Sec. 4.1),dnode [Dynamic Selection without Permanent Eviction (Sec 4.1.3), dnode [ InfLLM [121], Quest [122], PQCache [98], SqueezedAttention [123], RetrievalAttention [124], EM-LLM [125], ClsuterKV [126], wnode_1 ] ] [Dynamic Selection with Permanent Eviction (Sec 4.1.2), dnode [ H2O [127], BUZZ [128], NACL [129], Scissorhands [130], Keyformer [131], SepLLM [132],wnode_1 ] ] [Static KV Cache Selection (Sec 4.1.1), dnode [FastGen [133], SnapKV [134], Attention-Gate [135],wnode_1 ] ] ] ]

### v3

@@ Abstract
Large Language Models (LLMs) have revolutionized a wide range of domains such as natural language processing, computer vision, and multi-modal tasks due to their ability to comprehend context and perform logical reasoning. However, the computational and memory demands of LLMs, particularly during inference, pose significant challenges when scaling them to real-world, long-context, and real-time applications. Key-Value (KV) cache management has emerged as a critical optimization technique for accelerating LLM inference by reducing redundant computations and improving memory utilization. This survey provides a comprehensive overview of KV cache management strategies for LLM acceleration, categorizing them into token-level, model-level, and system-level optimizations. Token-level strategies include KV cache selection, budget allocation, merging, quantization, and low-rank decomposition, while model-level optimizations focus on architectural innovations and attention mechanisms to enhance KV reuse. System-level approaches address memory management, scheduling, and hardware-aware designs to improve efficiency across diverse computing environments. Additionally, the survey provides a comprehensive overview of both text and multi-modal datasets and benchmarks used to evaluate these various strategies. By presenting detailed taxonomies and comparative analyses, this work aims to offer valuable and actionable insights for researchers and practitioners to support the development of efficient and scalable KV cache management techniques, contributing to the practical deployment of LLMs in real-world applications. The curated paper list for KV cache management is in: https://github.com/TreeAI-Lab/Awesome-KV-Cache-Management.
@@ 1 Introduction
Large Language Models (LLMs) [1, 2, 3], trained on massive corpora, have revolutionized various domains such as natural language processing [4, 5, 6], computer vision [7, 8, 9], and multi-modal [10, 11, 12] tasks. Their ability to understand context and perform logical reasoning has enabled remarkable success in various fields, such as time series analysis [13, 14], recommendation [15, 16], autonomous driving [17, 18, 19], and healthcare [20, 21]. These breakthroughs are powered by state-of-the-art architectures and training paradigms, enabling models to achieve unparalleled performance across diverse tasks. Prominent LLMs, such as GPT [22, 23, 24], LLama [25, 26], DeepSeek [27, 28, 29], Mistral [30] [31], and GLM [32, 33], are built on the foundational transformer architecture [34], which excels at capturing long-range dependencies in sequential data. However, despite their powerful capabilities, the computational and memory demands of LLMs, particularly during inference, present significant challenges when scaling them to real-world, long-context, and real-time applications.
@@
A critical bottleneck in LLM inference lies in the efficient management of Key-Value (KV) pairs. Recently, caching techniques [35, 36] have been extensively employed to store previously computed intermediate results, allowing their reuse in subsequent inference steps to accelerate the model, such as graph neural networks [37, 38, 39]. Fortunately, the auto-regressive generation mechanism inherent to LLMs presents an opportunity to leverage KV caching for efficient text generation. Specifically, auto-regressive generation enables LLMs to produce text token by token, with each token conditioned on all previously generated ones. While this approach is highly effective for generating coherent and contextually relevant outputs, it struggles with poor scalability for long input sequences. This limitation arises because LLMs must compute attention values for every pair of tokens, causing the time and space complexity of the attention matrix to grow quadratically with sequence length. To address this issue, the KV cache mechanism stores the key and value matrices from previous decoding steps, allowing them to be reused. This significantly reduces redundant computations, as the model only computes attention values between new tokens and previously processed tokens, avoiding the need to recompute all attention values.
@@
Several recent surveys [3, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51] have explored the domain of efficient LLMs. These surveys primarily examine various aspects of LLM efficiency, presenting valuable insights while leaving room for further refinement and innovation. In particular, these existing works primarily focus on holistic approaches to improving LLM efficiency, systematically examining a wide range of techniques across multiple dimensions, such as data-level optimizations (e.g., prompt engineering), model architecture-level optimizations (e.g., efficient transformer designs), and system-level optimizations (e.g., task scheduling). For instance, Ding et al. [43] explore efficiency techniques that integrate data-level and model architecture perspectives, while Miao et al. [44] examine efficient LLM inference from a comprehensive system-level perspective. Similarly, Tang et al. [47], Wan et al. [45], and Xu et al. [49] provide analyses that encompass data, model, and system-level optimizations, reflecting holistic approaches to LLM acceleration.
@@
On the other hand, some surveys focus on more specialized aspects for LLM acceleration. For example, Zhu et al. [3], Park et al. [41], Wang et al. [42], and Tang et al. [47] focus on model compression as a key aspect of model-level optimization. Similarly, Kachris et al. [48] examines hardware acceleration strategies tailored for LLMs, while Xu et al. [49] investigates parameter-efficient tuning approaches. Albalak et al. [50] discusses data selection strategies to enhance the efficiency of LLM training, and Xia et al. [52] highlights collaborative techniques, such as speculative decoding [53, 54], to accelerate model inference. Li et al. [55] focuses on prompt compression. Similar to our work, Shi et al. [56], Li et al. [57], and Yuan et al. [58] also explore the use of KV caches to accelerate LLMs. However, our survey is both complementary and more comprehensive, offering a detailed taxonomy of KV cache management for text-based and multi-modal LLMs. We categorize techniques into token-level, model-level, and system-level perspectives and include benchmarks for both text and multi-modal scenarios. In particular, complementing existing KV cache surveys, we provide a detailed comparison of the differences and advantages of existing models at the token-level, model-level, and system-level.
@@ 2.1 Transformer Architecture
Transformers [34] have become the backbone of LLMs due to their ability to efficiently capture long-range dependencies in sequential data, such as text. This capability makes them particularly well-suited for tasks like machine translation, text generation, and image captioning. The transformer architecture follows an encoder-decoder structure, where most LLMs utilize only the decoder component. We first introduce the core components of the Transformer decoder and then describe the critical auto-regressive generation mechanism. Particularly, we do not describe certain components in the transformer, such as normalization, as they do not impact the understanding of KV cache management.
@@ 2.1.1 Transformer Decoder
As shown in Fig. 1, a decoder-based transformer architecture is composed of multiple stacked Transformer blocks, each designed to process sequential data effectively. Typically, a Transformer block consists of two core components, i.e., a Multi-Head Self-Attention (MHSA) mechanism and a Feed Forward Network (FFN). These blocks are arranged sequentially, where the output of one block is passed as input to the next. This iterative design allows the model to refine its understanding of the input sequence progressively, making it highly effective for tasks such as text generation and language modeling.
Positional Encoding. Before the input sequence is processed by the Transformer blocks, it undergoes a preprocessing phase. First, a tokenizer processes the input sentence $X$ by splitting it into discrete units, such as words or subwords. The resulting sequence can be represented as ${X}=[x_{1},x_{2},\cdots,x_{|X|}]$. These tokens are then mapped to dense vector representations using an embedding layer, i.e., $\mathbf{X}=\mathbf{I}_{X}\mathbf{E}^{\top}$, where $\mathbf{I}_{X}\in\{0,1\}^{n\times d_{\text{vocab}}}$ represents the one-hot vector of tokenized input $X$, $\mathbf{E}\in\mathbb{R}^{d_{\text{vocab}}\times d_{x}}$ is the embedding matrix, and $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{|X|}]\in\mathbb{R}^{n\times d_{x}}$ is the resulting matrix of embedded token representations. Since the Transformer architecture does not inherently account for the order of tokens in a sequence, positional encodings are added to the token embeddings $\mathbf{X}$ to incorporate positional information. This can be expressed as $\mathbf{X}=\mathbf{X}+PE(X)$, where $PE(X)\in\mathbb{R}^{n\times d_{x}}$ represents a function [60, 61, 62] (e.g., RoPE [62]) that generates positional embeddings for the input ${X}$. Note that relative positional embeddings, such as RoPE [62] (Rotary Positional Embedding), differ significantly from absolute positional encoding. RoPE introduces positional information at each layer of the model through rotational transformations.
@@
where $\mathbf{W}_{Q_{i}}\in\mathbb{R}^{d_{x}\times d_{k}}$, $\mathbf{W}_{K_{i}}\in\mathbb{R}^{d_{x}\times d_{k}}$ and $\mathbf{W}_{V_{i}}\in\mathbb{R}^{d_{x}\times d_{v}}$ are the learned weight parameters. Then, the self-attention operation is applied to each triple $(\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i})$, and obtains the output of the $i$-th attention head $\mathbf{Z}_{i}$ as follows:
@@
where $\sqrt{d_{k}}$ is a scaling factor to ensure numerical stability. To capture diverse relationships, multiple attention heads with $h$ heads are applied to $\mathbf{X}$ in parallel, and their outputs are concatenated with one transformation as follows:
@@
where Concat is the concatenation operation and $\mathbf{W}_{O}\in\mathbb{R}^{d_{v}\times d_{o}}$ are the trainable parameters.
Figure 1: The decoder-only Transformer for LLMs.
Following the self-attention mechanism, the output is passed through a Feed Forward Network (FFN). The FFN is a fully connected neural network that applies two linear transformations separated by a nonlinear activation function $\sigma(\cdot)$ (e.g., ReLU [63]) :
@@
where $\mathbf{W}_{1}\in\mathbb{R}^{d_{o}\times d_{1}}$ and $\mathbf{W}_{2}\in\mathbb{R}^{d_{1}\times d_{2}}$ are two parameters, $\mathbf{b}_{1}\in\mathbb{R}^{d_{1}}$ and $\mathbf{b}_{2}\in\mathbb{R}^{d_{2}}$ are two bias vectors.
@@ 2.2.1 Auto-regressive Generation with KV Cache
Formally, at each decoding step $t$, the new token embedding $\mathbf{x}_{t}$ is used to compute the query vector $\mathbf{q}^{t}_{i}$, key vector $\mathbf{k}^{t}_{i}$, and value vector $\mathbf{v}^{t}_{i}$ as follows:
@@
where $\mathbf{\hat{K}}_{i}^{t-1}\in\mathbb{R}^{t-1\times d_{k}}$ and $\mathbf{\hat{V}}_{i}^{t-1}\in\mathbb{R}^{t-1\times d_{v}}$ represent the cached key and value matrices of tokens in $X[1:t-1]$. These cached matrices are then used in the scaled dot-product attention computation for the token $x_{t}$. The attention output $\mathbf{z}^{t}_{i}$ for the token $x_{t}$ at step $t$ is calculated as:
@@ 2.2.2 Time and Space Complexity Analysis
Given a transformer-based $L$-layer LLM with $h$ attention heads per layer and an input sequence of length $X=[x_{1},\cdots,x_{t}]$, we analyze the time saved and the space required to store cached KV pairs. For simplicity, we assume that the keys and values of the $t_{c}$ tokens are stored for all heads across all LLM layers.
Computational Complexity. For each token, the saved computation time comes from avoiding the repeated computation of Keys and Values in Equation 1, the self-attention result in Equation 2, and the linear transformation in Equation 3. We omit the time analysis on operations in transformer that do not affect the understanding of KV cache acceleration, such as layer norm and position encoding.
  • QKV Computation. The time of computing Queries, Keys and Values for each token in Equation 1 is $\triangle_{1}=O(2d_{x}d_{k}+d_{x}d_{v})$.
  • Self-attention Result. The time complexity of computing each attention score $\mathbf{z}_{i}$ in Equation 2 takes $O(t(d_{k}+d_{v}))$.
  • Linear Transformation. To merge the $h$ attention results in Equation 3 the time is $\triangle_{2}=O(hd_{v}+d_{v}d_{o})$.
@@
Space Complexity. Compared to computation without caching, additional space is required to store the cached KV pairs for $t_{c}$ tokens across $h$ attention heads and $L$ layers. Assuming each Key and Value is stored in Float16 precision, the total extra space needed can be expressed as:
$O(L\cdot h\cdot t_{c}\cdot(d_{k}+d_{v})\cdot sizeof(Float16))$ (11)
@@ 2.3 Challenges in KV Cache Management
  • Cache Eviction Policies: Determining which items to evict when the cache reaches its capacity is a complex problem. Popular policies [36] like Least Recently Used (LRU) or Least Frequently Used (LFU) do not align with LLM patterns, leading to suboptimal performance.
@@ 3 Taxonomy
In the above sections, we analyzed how the number of cached Key-Value (KV) pairs significantly impacts both the computation time and the additional memory required during inference. Efficient KV cache management is critical to balancing performance improvements and resource utilization, especially as sequence lengths and model sizes continue to grow. After carefully reviewing existing approaches, we categorize KV cache optimization strategies into three levels: token-level optimization, model-level optimization, and system-level optimization. Each level addresses specific aspects of the challenges associated with KV cache management and offers distinct techniques to enhance efficiency. The detailed taxonomy is illustrated in Fig. 2.
  • Token-Level Optimization refers to improving KV cache management efficiency by focusing on the fine-grained selection, organization, and compression at the token level, requiring no architectural changes to the original model. While KV cache selection (Sec. 4.1) focuses on prioritizing and storing only the most relevant tokens, KV cache budget allocation (Sec. 4.2) dynamically distributes memory resources across tokens to ensure efficient cache utilization under limited memory. Furthermore, KV cache merging (Sec. 4.3) reduces redundancy by combining similar or overlapping KV pairs, while KV Cache Quantization (Sec. 4.4) minimizes the memory footprint by reducing the precision of cached KV pairs. Finally, KV cache low-rank decomposition (Sec. 4.5) uses low-rank decomposition techniques to reduce cache size.
  • Model-level Optimization refers to designing an efficient model structure to optimize KV cache management. This can further refer to several strategies: Attention grouping and sharing (Sec. 5.1) methods examine the redundant functionality of keys and values and group and share KV cache within or across transformer layers. Architecture alterations (Sec. 5.2) emerge to design new attention mechanisms or construct extrinsic modules for KV optimization. Furthermore, there are also works designing or combining non-transformer architectures (Sec. 5.3) that adopt other memory-efficient designs like recurrent neural networks to optimize the KV cache in traditional transformers.
@@
Decomposition (Sec 4.5.2), dnode [ DecoQuant [66] , wnode_1 ] ] [Singular Value Decomposition (Sec 4.5.1), dnode [ ECKVH [67], EigenAttention [68], ZDC [69], LoRC [70], ShadowKV [71], Palu [72] Q-Filters [73], Loki [74], , wnode_1 ] ] ] [KV Cache
@@
Merging (Sec 4.3.1), dnode [ CCM [104], LoMA [105], DMC [106], CaM [107], D2O [108], AIM [109], Look-M [110], ZeroMerge [111], KVMerger [112], CHAI [113], wnode_1 ] ] ] [KV Cache Budget Allocation (Sec. 4.2), dnode [Head-wise Budget Allocation (Sec 4.2.2), dnode [ AdaKV [114], CriticalKV [115], LeanKV [116], RazorAttention [117], HeadKV [118], DuoAttention [119] ,wnode_1 ] ] [Layer-wise Budget Allocation (Sec 4.2.1), dnode [ PyramidKV [120], PyramidInfer [121], DynamicKV [122], PrefixKV [123], SimLayerKV [124] , wnode_1 ] ] ] [KV Cache
Selection (Sec. 4.1),dnode [Dynamic Selection without Permanent Eviction (Sec 4.1.3), dnode [ LoopServe [1], InfLLM [125], Quest [126], PQCache [101], SqueezedAttention [127], RetrievalAttention [128], EM-LLM [129], ClsuterKV [130], Loki [74], wnode_1 ] ] [Dynamic Selection with Permanent Eviction (Sec 4.1.2), dnode [ H2O [131], BUZZ [132], NACL [133], Scissorhands [134], Keyformer [135], SepLLM [136],wnode_1 ] ] [Static KV Cache Selection (Sec 4.1.1), dnode [FastGen [137], SnapKV [138], L2Compress [139], Attention-Gate [140],wnode_1 ] ] ] ]